# Learning in Structured Stackelberg Games

### v1

@@ 4.2.2 Stackelberg ERM learner
Consider a hypothesis class $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$, a Stackelberg game, $\pazocal{G}$, and let $S\subseteq\pazocal{Z}$. We say that $\pazocal{H}$ SN-shatters $S$ with respect to game $\pazocal{G}$, if there exist $\phi_{1},\phi_{2}:S\to\Delta(\pazocal{A})$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that:
@@
Example 4.11.
Consider the hypothesis class of conjunctions, where each hypothesis is defined over a set of Boolean variables, $\pazocal{Z}=\{z_{1},\cdots,z_{n}\}$, such that $\pazocal{H}_{\text{conj}}=\{h:\{0,1\}^{n}\to\{0,1\}|h(z)=\land_{i\in S}z_{i}\land_{j\in T}z_{j},S,T\subseteq[n],S\cap T=\emptyset\}$. Since each hypothesis outputs in binary, this hypothesis class could naturally be considered in a Stackelberg game with two follower types, $f^{(0)},f^{(1)}$. The Natarajan dimension in this case reduces to the well-studied notion of the VC-dimension, which is known to be equal to $n$ for $\pazocal{H}_{\text{conj}}$. At the same time, the $d_{SN,\pazocal{G}}(\pazocal{H}_{\text{conj}})$ will be upper bounded by the size of the set $\{z\in\{0,1\}^{n}|\pi_{*}^{(0)}(z)\neq\pi_{*}^{(1)}(z)\}$, which could be as small as $1$ for two distinct follower types.
@@
Consider the Stackelberg ERM policy, $\pi_{\text{S-ERM}}$, as in Definition 4.9 and let $h_{\text{S-ERM}}$ be the corresponding hypothesis in $\pazocal{H}$. Then universally for any such policy, there exists constant $C$, such that for every $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ and distribution $\mathscr{D}$,
Proof Sketch: The proof follows from a similar argument as that of the sample complexity bounds of multiclass learning [17]. The only part of the proof that does not apply here is Natarajan’s lemma. We prove an appropriate generalization in Lemma 4.13. The full proof of Theorem 4.12 is presented in Appendix B.
Lemma 4.13.
$|\Pi_{\pazocal{H},\pazocal{G}}|\leq|\pazocal{Z}|^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$.
We present the complete proof of Lemma 4.13 in Appendix B.
Remark 4.14.
It is worth noting that even when there is only one context (i.e. $\pazocal{Z}$ is a singleton), our results in this section are incomparable with those in Section 4.1. This is because in this section, the goal of the leader is to learn a deterministic mapping from contexts to follower types, assuming that one exists. So, if the contexts are the same between all examples in the training set, the follower types corresponding to these examples will also be the same. In contrast, the context-free setting we consider in Section 4.1 allows for a distribution over follower types. Thus the two bounds are generally incomparable.
5 Conclusion
In this work we study structured Stackelberg games, where contextual information is predictive of the follower’s type. The follower type that the leader is playing against is unknown, while the context that carries information about the state of the world is presented to both players. We show that the leader can use the unknown, fixed mapping from contexts to follower types to inform her strategy. In the online setting, our algorithms learn the optimal policy for the leader and achieve tight no-regret bounds. These bounds scale with a new notion of statistical complexity that directly exploits the structure of Stackelberg games. We show similar results in the distributional setting, where contexts and follower types are drawn from a (joint) distribution.
An interesting direction for future research is to analyze structured Stackelberg games in the beyond worst-case analysis framework. It has recently been shown that smooth online learning is as easy as distributional learning, with regret bounds scaling with the VC dimension of the hypothesis class. Extending these findings to multiclass learning and structured Stackelberg games might lead to improved results both from a statistical and computational point of view.
Appendix A Details and Proofs from Section 3
@@ A.1 A technical subtlety
An important component of our analysis is that the leader must be able to compute the optimal strategy for a given context and follower type. For this, we state that for any context $\mathbf{z}\in\pazocal{Z}$ the leader incurs negligible regret by restricting herself to a subset $\mathscr{E}_{\mathbf{z}}\subseteq\Delta(\pazocal{A})$ which is finite and computable. To formally state our result we present the notion of a contextual best response region, which was first introduced in [24] and which we repeat here for completeness.
@@
Fix a context $\mathbf{z}\in\pazocal{Z}$ and consider the set of all non-empty contextual best-response regions. For $\delta>0$, $\mathscr{E}_{\mathbf{z}}(\delta)$ is the set of leader mixed strategies such that for all best-response functions $\sigma$ and any $\mathbf{x}\in\Delta(\pazocal{A})$ that is an extreme point of cl$(\pazocal{X}_{\mathbf{z}}(\sigma))$, $\mathbf{x}\in\mathscr{E}_{\mathbf{z}}(\delta)$ if $\mathbf{x}\in\pazocal{X}_{\mathbf{z}}(\sigma)$. Otherwise there is some $\mathbf{x}^{\prime}\in\mathscr{E}_{\mathbf{z}}(\delta)$ such that $\mathbf{x}^{\prime}\in\pazocal{X}_{\mathbf{z}}(\sigma)$ and $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{1}\leq\delta$.

### v2

@@ 4.4 Upper bound
Let $\pazocal{H}:[K]^{\pazocal{Z}}$ and $\pazocal{G}$ be a Stackelberg game. We say that a sample $S\in\pazocal{Z}^{n}$ is shattered by $\pazocal{H}$ with respect to $\pazocal{G}$ if there exists a function $g:\pazocal{Z}\to[K]$ such that for all $b\in\{0,1\}^{n}$ there exists an $h_{b}\in\pazocal{H}$ such that:
  • $h_{b}(\mathbf{z}_{i})=g(i)\;\;\forall i\in[n]$ such that $b_{i}=0$, and
  • $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g(\mathbf{z}_{i}),h_{b}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})\geq\gamma\;\;\forall i\in[n]$ such that $b_{i}=1$.
Definition 4.13 ($\gamma$-valued Stackelberg-Graph Dimension).
The $\gamma$-valued Stackelberg-Graph dimension, denoted as $\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$ is the cardinality of the largest set that is $\gamma$-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
We are not ready to show our main result in this section, which shows that the $\gamma$-Graph dimension upper bounds the PAC sample complexity in our setting. We design the following simple, improper algorithm that achieves the corresponding bound.
Algorithm 2 Stackelberg Distributional Learner ($\mathfrak{L}^{*}$)
1:input: sample $S=\{(\mathbf{z}_{1},f_{1}),\dots,(\mathbf{z}_{n},f_{n})\}$ that is realizable by $\pazocal{H}$ and test point $\mathbf{z}\in\pazocal{Z}$
2:Consider the subset of hypotheses consistent with the sample, $\pazocal{H}_{c}=\pazocal{H}|_{S}$.
3:Consider the set of followers predicted by $\pazocal{H}_{c}$ on test point $\mathbf{z}$, $F=\pazocal{H}_{c}(\mathbf{z})$
4:Compute and return $\mathbf{x}^{*}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in F}r(\mathbf{z},\mathbf{x},f^{(i)})$
5:end
@@
For any class $\pazocal{H}:[K]^{\pazocal{Z}}$, any Stackelberg game, $\pazocal{G}$, and for any $\epsilon,\delta,\gamma\in(0,1)$, there exists some constant $C_{1}$ such that Algorithm˜2 achieves:
$m_{\mathfrak{L}^{*}}^{r}(\pazocal{H},\epsilon,\delta,\gamma)\leq C_{1}\frac{\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(1/\epsilon)+\log(1/\delta)}{\epsilon}.$
@@
We work with the cut-off loss problem for some parameters $(\epsilon,\delta,\gamma)\in(0,1)^{3}$. We define the cut-off loss
$\text{er}_{\mathscr{D},\gamma}(h):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\right]$
and the empirical cut-off loss for a dataset $S\in(\pazocal{Z}\times[K])^{n}$
$\hat{\text{er}}_{S,\gamma}(h):=\frac{1}{n}\sum_{(\mathbf{z}_{i},f_{i})\in S}\mathbbm{1}\biggl{\{}\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\biggr{\}}$
and the empirical loss
$\hat{\text{er}}_{S}(h):=\frac{1}{n}\sum_{(\mathbf{z}_{i},f_{i})\in S}\mathbbm{1}\biggl{\{}h(\mathbf{z}_{i})\neq h^{\prime}(\mathbf{z}_{i})\biggr{\}}.$
For $\pazocal{M}=\pazocal{Z}\times[K]$ we are interested in the following set
$Q_{\epsilon,\gamma}=\{M\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{M}(h)=0,\text{er}_{\mathscr{D},\gamma}(h)>\epsilon\}$
Now using a classic symmetrization argument [35, 3], we can upper bound the probability of $Q_{\epsilon,\gamma}$ with that of the following set
$R_{\epsilon,\gamma}=\{(N,S)\in\pazocal{M}^{n}\times\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{N}(h)=0,\hat{\text{er}}_{S,\gamma}(h)\geq\epsilon/2\}.$
The idea here is to bound the probability that there exists a hypothesis yielding maximum utility for the leader over the sample, but is $\gamma$-far from the optimal hypothesis, by the probability that this hypothesis has poor performance on another sample. More concretely, it can be shown that $\mathscr{D}^{n}(Q_{\epsilon,\gamma})\leq 2\mathscr{D}^{2n}(R_{\epsilon,\gamma})$ for $n\geq c/\epsilon$, for some constant $c$. To see this, we first write
$\mathscr{D}^{2n}(R_{\epsilon,\gamma})=\int_{Q_{\epsilon,\gamma}}\mathscr{D}^{n}(\{S\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{N}(h)=0,\hat{\text{er}}_{S,\gamma}(h)\geq\epsilon/2\})d\mathscr{D}^{n}(N)$
For some $N\in Q_{\epsilon,\gamma}$ consider some $h_{N}\in\pazocal{H}$ that satisfies $\hat{\text{er}}_{N,\gamma}(h_{N})=0$ and $\text{er}_{\mathscr{D},\gamma}(h_{N})>\epsilon$. We show that $\mathscr{D}(\hat{\text{er}}_{S,\gamma}(h_{S})\geq\epsilon/2)\geq 1/2$, from which the claim follows straightforwardly. We know that $\text{er}_{\mathscr{D},\gamma}(h_{N})>\epsilon$, and therefore $n\cdot\hat{\text{er}}_{S,\gamma}(h_{N})\leq n\cdot\hat{\text{er}}_{S}(h_{N})$ follows a binomial distribution with probability of success at least $\epsilon$. We note that the inequality above follows since mispredicting the follower type is a necessary condition for suboptimal utility. Using a Chernoff bound, we can show that
$\mathscr{D}^{n}(S:\hat{\text{er}}_{S,\gamma}(h_{N})<\epsilon/2)\leq e^{-n\epsilon/8}$
and for $n=c/\epsilon$, we get the desired result.
The next step is to use a random swap argument to upper bound $\mathscr{D}^{2n}(R_{\epsilon,\gamma})$ with something that is independent of $\mathscr{D}$. We let $\Gamma_{n}$ denote the set of all permutations on $\{1,\dots,2n\}$ that potentially swap the $i$’th and $n+i$’th elements for all $i\in[n]$. So for all $\sigma\in\Gamma_{n}$ and $i\in[n]$, we have that either $\sigma(i)=i,\sigma(n+i)=n+i$ or $\sigma(i)=n+i,\sigma(n+i)=i$. From Lemma 3 in Attias et al. [3] (cf. also [2]), we get that for some $\pazocal{M}=\pazocal{Z}\times[K]$ and distribution $\mathscr{D}$ on $V$:
$\mathscr{D}^{2n}(R_{\epsilon,\gamma})=\mathop{\mathbb{E}}_{Z\sim\mathscr{D}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}]\leq\max_{Z\in\pazocal{Z}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}].$
The last step of the proof, requires reducing to learning a partial binary hypothesis class and using a disambiguation argument to turn the partial class into a total one, as in Attias et al. [3]. Concretely, we start by fixing a set $V\in\pazocal{M}^{2n}$, where $v_{i}=(\mathbf{z}_{i},h^{*}(\mathbf{z}_{i}))$ and a set $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{2n}\}$. For the projection, $\pazocal{H}(S)$ of the hypothesis class on the set, one can define the partial binary hypothesis class
$\pazocal{H}:=\left\{h^{\prime}\in\{0,1,\star\}^{2n}:\exists h\in\pazocal{H}|_{S}:\forall i\in[2n]\;\;h^{\prime}(\mathbf{z}_{i})=\begin{cases}0\;\;\text{if}\;\;h(\mathbf{z}_{i})=h^{*}(\mathbf{z}_{i})\\ 1\;\;\text{if}\;\;\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\\ \star\;\;\text{if}\;\;\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})\leq\gamma\end{cases}\right\}$
It is easy to verify that $\textnormal{VCdim}(\pazocal{H})\leq\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$. We can now turn this partial concept class into a total concept class of size $n^{O(\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(n))}$ using the following disambiguation argument.
Lemma 4.15 ([1]).
Let $\pazocal{H}^{\prime}$ be a partial concept class on a finite space $S$, with $\textnormal{VCdim}(\pazocal{H}^{\prime})=d$. Then there exists a disambiguation $\tilde{\pazocal{H}}^{\prime}$ of $\pazocal{H}^{\prime}$ of size $|\tilde{\pazocal{H}}^{\prime}|=|S|^{O(d\log|S|)}$.
We have that $\sigma(V)\in R_{\epsilon,\gamma}$ iff there exists some $h\in\pazocal{H}$ such that $h(\mathbf{z}_{\sigma(i)})=h^{*}(\mathbf{z}_{\sigma(i)})\;\;\;\forall\;i\in[n]$ and
$\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\biggl{\{}\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}_{\sigma(n+i)}),h^{*}(\mathbf{z}_{\sigma(n+i)})\}}r(\mathbf{z}_{\sigma(n+i)},\mathbf{x},f^{(j)})>\gamma\biggr{\}}\geq\epsilon/2.$
We know that if there exists such an $h$ that is correct on the first $n$ points but $\gamma$-far on at least $\epsilon n/2$ of the last $n$ points, there exists a function in $\tilde{\pazocal{H}}^{\prime}$, that is 0 on the first $n$ points and 111 on the points on which $h$ is $\gamma$-far. We also know that if for some $\sigma\in\Gamma_{n}$, $\tilde{h}^{\prime}\in\pazocal{H}^{\prime}$ is a witness that $\sigma(V)\in R_{\epsilon,\gamma}$, then at least one of $\tilde{h}^{\prime}(i),\tilde{h}^{\prime}(i)$ must be 0, and at least $n\epsilon/2$ elements must be non-zero. Thus for a permutation drawn uniformly at random, the probability that all 1’s are on the last n elements is at most $2^{-n\epsilon/2}$. Also since $|\tilde{\pazocal{H}}^{\prime}|$ is bounded we an use a union bound that yields $\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}]\leq n^{O(\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(n))}2^{-n\epsilon/2}$. Given that $V\in\pazocal{M}^{2n}$, we can take a $\max$ over all $V\in\pazocal{M}^{2n}$, which gives the desired result. ∎
Remark 4.16.
It is worth noting that even when there is only one context (i.e. $\pazocal{Z}$ is a singleton), our results in this section are incomparable with those in Section˜4.1. This is because in this section, the goal of the leader is to learn given a deterministic mapping from contexts to follower types. So, if the contexts are the same between all examples in the training set, the follower types corresponding to these examples will also be the same. In contrast, the context-free setting we consider in Section˜4.1 allows for a distribution over follower types. Thus the two bounds are generally incomparable.
5 Computational Complexity Considerations
Computational complexity has been a longstanding concern in the analysis of Stackelberg games. In the simplest setting, where a leader interacts with a single known follower type Conitzer and Sandholm [20] have shown that computing the optimal mixed strategy for the leader can be achieved in polynomial time using linear programming. However, in more complex Stackelberg games such as Bayesian Stackelberg games, Stackelberg security games, and even 3-player Stackelberg games computing the leader’s optimal strategy has been shown to be NP-hard [20, 28]. Therefore computational considerations have so far been largely ignored. In our setting, there are several interesting cases of structured Stackelberg games for which the leader’s optimal strategy can be learned in polynomial time with provable (albeit not optimal) regret guarantees.
Observation.
In a structured Stackelberg game with hypothesis class $\pazocal{H}$ the leader’s optimal policy can be learned in polynomial time if $\pazocal{H}$ is efficiently online learnable. We give a high-level overview of this approach. First, we consider a learning algorithm for the hypothesis class $\pazocal{H}$. The leader can use this algorithm to predict the follower’s type and based on that compute the optimal mixed strategy against that type as if her prediction were correct. As shown by Conitzer and Sandholm [20], for a fixed and known follower type, computing the leader’s optimal mixed strategy can be achieved in polynomial time using linear programming. The learner will make a finite number of follower type mispredictions before finding $h^{*}$, and at this point will play optimally against every follower type that appears in every subsequent round. Therefore, for any hypothesis class in the realizable setting that is efficiently online- (and therefore also PAC-) learnable we can establish a computationally efficient algorithm for learning the leader’s optimal policy. We now present one such example of a hypothesis class that can be learned efficiently when used to predict the followers’ types.
Example.
[…]
@@ 5 Computational Complexity Considerations
The class of multiclass decision lists is efficiently online learnable.
We present the full proof of Theorem˜5.1 and the corresponding algorithm in Section˜A.5.
6 Conclusion and Future Directions
We initiate the study of structured Stackelberg games, where contextual information is predictive of the follower’s type. Our main results include a complete characterization of instance-optimal regret bounds in the online setting and PAC sample complexity guarantees in the distributional setting. In this section, we discuss concrete open problems and future directions.
Computationally efficient algorithms.
An interesting avenue for future research is to consider the computational complexity of structured Stackelberg game algorithms. Although we have shown how to achieve optimal regret bounds, our proposed algorithm is not computationally efficient. As we discuss in Section˜5, our model opens the door for efficient, yet so far sub-optimal learning algorithms in certain settings. It would be interesting to further explore whether the structure introduced in our work can lead to both optimal and efficient algorithms or if there is an underlying tradeoff between computational complexity and tight regret guarantees.
Regret guarantees for the agnostic setting.
Our work focuses on structured Stackelberg games in the realizable setting, where the leader is given a class of hypotheses with the promise that it includes the optimal predictor. The agnostic learning setting is a natural relaxation of this assumption. While we provide some results (cf. Section˜B.1.4) for this setting, our bounds are not instance optimal. Providing a complete characterization for the agnostic case, analogously to the realizable one we study here, remains an avenue for future research. It is plausible that an entirely new complexity measure will be required here, which would show a clear separation between our setting and multiclass classification, where the Littlestone dimension characterizes both the realizable and the agnostic setting.
Beyond worst-case analysis frameworks.
Another direction for future research is to analyze structured Stackelberg games in the beyond worst-case analysis framework. It has recently been shown that smooth online learning is as easy as distributional learning, with regret bounds scaling with the VC dimension of the hypothesis class. Extending these findings to multiclass learning and structured Stackelberg games might lead to improved results both from a statistical and computational point of view.
Appendix A Details and Omitted Proofs
@@ A.1 A technical subtlety
An important component of our analysis is that the leader must be able to compute the optimal strategy for a given context and follower type. For this, we state that for any context $\mathbf{z}\in\pazocal{Z}$ the leader incurs negligible regret by restricting herself to a subset $\mathscr{E}_{\mathbf{z}}\subseteq\Delta(\pazocal{A})$ which is finite and computable. To formally state our result we present the notion of a contextual best-response region, which was first introduced in [27] and which we repeat here for completeness.
@@
[…]